# From Discrete to Continuous Binary Best-Response Dynamics: Discrete Fluctuations Almost Surely Vanish with Population Size

### v2

@@ V-D Global stability analysis
The abstract dynamics and in turn the semicontinuous population dynamics may admit several clean-cut, anticoordinator-driven, or coordinator-driven equilibrium points. Denote the set of clean-cut, anticoordinator-driven, and coordinator-driven equilibrium points of the semicontinuous population dynamics by $\bm{\mathcal{Q}}^{c}$, $\bm{\mathcal{Q}}^{a}$, and $\bm{\mathcal{Q}}^{o}$, respectively. When there is no equilibrium point of either type, the corresponding set will be empty. Assume that there are altogether $\mathtt{Q}\in\mathbb{N}$ equilibrium points of clean-cut or anticoordinator-driven type. Arrange the clean-cut or anticoordinator-driven equilibrium points of the semicontinuous population dynamics (5) in the ascending order of their associated equilibrium points in the abstract dynamics. Hence, we have
@@
[…]
According to 1, in the abstract dynamics, between every two anticoordinator-driven and/or clean-cut equilibria $q^{*}_{k}$ and $q^{*}_{k+1}$, $k\in[\mathtt{Q}-1]$, there exists exactly one coordinator-driven equilibrium, which we denote by $q^{*}_{k,k+1}$. The associated equilibrium in the semicontinuous dynamics is denoted by $\mathbf{q}^{*}_{k,k+1}$. Define $q^{*}_{0,1}=0$ and $q^{*}_{\mathtt{Q},\mathtt{Q}+1}=1$.
@@
Consider the semicontinuous population dynamics (5) and its associated abstract dynamics (9). Under 1,
@@
    The limit set of every point in $\{\bm{x}\in\bm{\mathcal{X}}_{ss}|x={q}^{*}_{k,k+1}\}$ for $k\in[\mathtt{Q}-1]$ is ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\{}\bm{q}^{*}_{k},\bm{q}^{*}_{k+1},\bm{q}^{*}_{k,k+1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\}}$.
@@
Under 1, for the semicontinuous population dynamics (5) and the associated abstract dynamics (9),
@@
V-E The Birkhoff center of the semicontinuous population dynamics
@@
If Assumption 1 holds, the Birkhoff center of the dynamical system induced by the semicontinuous population dynamics (5) is
@@
The following corollaries are direct results of Proposition 3 and determine the Birkhoff center of the dynamical system induced by the semicontinuous population dynamics capturing the behaviour of a population of exclusively coordinators and exclusively anticoordinators, respectively.
@@
If Assumption 1 holds, the Birkhoff center of the semicontinuous population dynamics of a population of exclusively coordinators is
@@
If Assumption 1 holds, the Birkhoff center of the semicontinuous population dynamics describing the evolution of an exclusive population of $\mathsf{p}$ types of anticoordinators is
@@
So far, the steady-state behaviour of the semicontinuous population dynamics (5) was determined. The following section determines the asymptotic behaviour of the discrete population dynamics (3) as the population size approaches infinity.
VI The asymptotic behaviour of the discrete population dynamics
@@
In words, consider a population of coordinators and anticoordinators who update their decisions asynchronously based on (1) and (2), respectively. As the population size approaches infinity, the population state almost surely only visits the states close to the equilibria. Consequently, the length of the possible perpetual fluctuations in the population proportion of $\mathtt{A}$-players will converge to zero. This results in the following corollary.
@@
Consider the population dynamics (3). Under 1, and starting from a specified initial condition, as the population size approaches infinity, the length of the fluctuations in the population proportion of $\mathtt{A}$-players converges to zero with probability one.
Nevertheless, we can still have perpetual fluctuations (with nonzero probability), but the length of the fluctuations will vanish with the population size almost surely. Roughly speaking, the state if does not equilibrate will reach a non-singleton minimally positively invariant set, but for “large” populations most of the time in the long run will spend “close” to the isolated members of the Birkhoff center.

### v3

@@ V-D Global stability analysis
The abstract dynamics and, in turn, the continuous-time population dynamics may admit several clean-cut, anticoordinator-driven, or coordinator-driven equilibrium points. Denote the set of clean-cut, anticoordinator-driven, and coordinator-driven equilibrium points of the continuous-time population dynamics by $\bm{\mathcal{Q}}^{c}$, $\bm{\mathcal{Q}}^{a}$, and $\bm{\mathcal{Q}}^{o}$, respectively. When there is no equilibrium point of either type, the corresponding set will be empty. Assume that there are altogether $\mathtt{Q}\in\mathbb{N}$ equilibrium points of clean-cut or anticoordinator-driven type. Arrange the clean-cut or anticoordinator-driven equilibrium points of the continuous-time population dynamics (7) in the ascending order of their associated abstract equilibrium points. Hence, we have
@@
where $q^{*}_{k}$ is the abstract state at $\mathbf{q}^{*}_{k}\in\bm{\mathcal{Q}}^{a}\cup\bm{\mathcal{Q}}^{c}$ for $k\in[\mathtt{Q}]$. A set $\bm{\mathcal{M}}$ is attractive under differential inclusion (4) from a set $\bm{\mathcal{U}}$ if for each solution $\bm{x}(t)$ with $\bm{x}(0)\in\bm{\mathcal{U}}$ and each open $\epsilon$-neighborhood of $\bm{\mathcal{M}}$, there exists time $T>0$ such that $\bm{x}(t)$ falls in $\epsilon$-neighborhood of $\bm{\mathcal{M}}$ for all $t\geq T$ [45]. The union of all sets from which $\bm{\mathcal{M}}$ is attractive under differential inclusion $\bm{\mathcal{V}}$ is the basin of attraction of $\bm{\mathcal{M}}$ under $\bm{\mathcal{V}}$ [45].
According to A2, in the abstract dynamics, between every two anticoordinator-driven and/or clean-cut equilibria $q^{*}_{k}$ and $q^{*}_{k+1}$, $k\in[\mathtt{Q}-1]$, there exists exactly one coordinator-driven equilibrium, which we denote by $q^{*}_{k,k+1}$. The associated equilibrium in the continuous-time dynamics is denoted by $\mathbf{q}^{*}_{k,k+1}$. Define $q^{*}_{0,1}=0$ and $q^{*}_{\mathtt{Q},\mathtt{Q}+1}=1$. The following theorem reveals the global behavior of the continuous-time population dynamics.
@@
Consider the continuous-time population dynamics (7) and its associated abstract dynamics (11). Under 1,
@@
    The limit set of every point in $\{\bm{x}\in\bm{\mathcal{X}}_{ss}|x={q}^{*}_{k,k+1}\}$ for $k\in[\mathtt{Q}-1]$ is ${\{}\bm{q}^{*}_{k},\bm{q}^{*}_{k+1},\bm{q}^{*}_{k,k+1}{\}}$.
@@
Under 1, for the continuous-time population dynamics (7) and the associated abstract dynamics (11),
@@
V-E The Birkhoff center of the continuous-time population dynamics
@@
If Assumption 1 holds, the Birkhoff center of the dynamical system induced by the continuous-time population dynamics (7) is
@@
The following corollaries are direct results of Proposition 3 and determine the Birkhoff center of the dynamical system induced by the continuous-time population dynamics capturing the behavior of a population of exclusively coordinators and exclusively anticoordinators, respectively.
@@
If Assumption 1 holds, the Birkhoff center of the continuous-time population dynamics of a population of exclusively coordinators is
@@
If Assumption 1 holds, the Birkhoff center of the continuous-time population dynamics describing the evolution of an exclusive population of $\mathsf{p}$ types of anticoordinators is
@@
So far, the steady-state behavior of the continuous-time population dynamics (7) was determined. The following section determines the asymptotic behavior of the discrete population dynamics (3) as the population size approaches infinity.
VI The asymptotic behavior of the discrete population dynamics
@@
In words, consider a population of coordinators and anticoordinators who update their decisions asynchronously based on (1) and (2), respectively. As the population size approaches infinity, the population state almost surely only visits the states close to the equilibria. Consequently, the amplitude of the possible perpetual fluctuations in the population proportion of $\mathtt{A}$-players will converge to zero. This results in the following corollary.
@@
Consider the population dynamics (3). Under 1 and starting from a given initial condition, as the population size approaches infinity, the amplitude of the fluctuations in the population proportion of $\mathtt{A}$-players converges to zero with probability one.
Nevertheless, we can still have perpetual fluctuations (with nonzero probability), but the amplitude of the fluctuations will vanish with the population size almost surely. Roughly speaking, the state if does not equilibrate for “large” populations most of the time in the long run will spend “close” to the isolated members of the Birkhoff center.
@@
1 ensures that $\bm{\mathcal{Q}}^{c}$, $\bm{\mathcal{Q}}^{a}$, and $\bm{\mathcal{Q}}^{o}$ are distinct. Otherwise, the following possible cases may happen: (i) The threshold of benchmark anticoordinating subpopulation $i$ equals ${\eta}^{\prime}_{j}+{\eta}_{i-1}$ (resp. ${\eta}^{\prime}_{j}+{\eta}_{i}$). In this case, following the approach taken in the proofs of Lemma 2 and Theorem 2, it can be shown that the equilibria $\mathbf{a}^{ij}$ and $\mathbf{c}^{(i-1)j}$ (resp. $\mathbf{c}^{ij}$) match and remain stable. (ii) The threshold of benchmark coordinating subpopulation $j$ equals ${\eta}^{\prime}_{j-1}+{\eta}_{i}$ (resp. ${\eta}^{\prime}_{j}+{\eta}_{i}$). In this case, following the approach taken in the proofs of Lemma 2 and Theorem 2, it can be shown that the equilibria $\mathbf{o}^{ij}$ and $\mathbf{c}^{i(j-1)}$ (resp. $\mathbf{c}^{ij}$) match, resulting in an unstable equilibrium. (iii) The threshold of benchmark coordinating subpopulation $j$ equals that of benchmark anticoordinating subpopulation $i$. In this case, the point $x=\tau_{i}$ is an equilibrium for the abstract dynamics if $\tau_{i}\in[{\eta}_{i-1}+{\eta}_{j-1}^{\prime},{\eta}_{i}+{\eta}_{j}^{\prime}]$. Following the approach taken in the proof of Theorem 2, it can be shown that if $\tau_{i}\in({\eta}_{i-1}+{\eta}_{j}^{\prime},{\eta}_{i}+{\eta}_{j-1}^{\prime})$, it is asymptotically stable, and unstable otherwise. In this case a continuum of equilibria $\mathbf{z}^{ij}$ emerges for the semicontinouous population dynamics. This continumm is characterized by $\bm{\mathcal{Z}}=\{\bm{x}\in\bm{\mathcal{X}}_{ss}|\forall l\notin\{i,\mathsf{p}+\mathsf{p}^{\prime}-j+1\},z^{ij}_{l}=a^{ij}_{l},z^{ij}_{i}+z^{ij}_{\mathsf{p}+\mathsf{p}^{\prime}-j+1}=\tau_{i}-({\eta}_{i-1}+{{\eta}_{j-1}^{\prime}})\}$. In words, all members of those coordinating subpopulations with thresholds less (resp. greater) than $\tau_{j}^{\prime}$ play strategy $\mathtt{A}$ (resp. $\mathtt{B}$), and all members of those anticoordinating subpopulations with thresholds less (resp. greater) than $\tau_{j}$ play strategy $\mathtt{B}$ (resp. $\mathtt{A}$). The proportions of strategy $\mathtt{A}$ players in the two benchmark subpopulations add up to $\tau_{i}-({\eta}_{i-1}+{{\eta}_{j-1}^{\prime}})$. The Birkhoff center will then also contain $\bm{\mathcal{Z}}$. The population proportion of strategy-$\mathtt{A}$ players is constant at $\bm{\mathcal{Z}}$ and consequently, the amplitude of the fluctuations in the population proportion of strategy-$\mathtt{A}$ players in the discrete population dynamics vanished with probability one.
Remark 5.
In some setups, such as in [15], agents include themselves when determining their preferred strategy. As detailed in Section VII-D, our main results hold for such update rules.
Revisiting Example 2.
Consider the associated continuous-time population dynamics with the finite population dynamics introduced in Example 2. Based on Lemma 2, the mean dynamics admit only one equilibrium point at $x=\tau_{1}$, which is an anticoordinator-driven equilibrium point. Therefore, according to 1 this equilibrium point is globally asymptotically stable. Figure 3 depicts the evolution of population proportion of strategy-$\mathtt{A}$ players merged with Figure 1. As seen, $x(t)$ approaches $\tau_{1}$. The chattering observed in the solid black line is due to the numerical errors.
Example 4.
Consider a population of size $\mathsf{N}$ consisting of four coordinating and three anticoordinating subpopulations. The population distribution over the subpopulations is $\bm{\rho}=({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2/28},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}3/28},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}3/28},{6/28},{8/28},{3/28},{3/28})$. The thresholds of the anticoordinating subpopulations are ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tau_{1}}={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0.929}$, ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tau_{2}}={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}6/7}$, ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tau_{3}}={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0.357}$, and those of coordinators are ${\tau^{\prime}_{1}}={0.05}$, ${\tau^{\prime}_{2}}={0.321}$, ${\tau^{\prime}_{3}}={0.5}$, ${\tau^{\prime}_{4}}={0.643}$. When the population size $\mathsf{N}$ equals $28$ or $56$ (which is the next larger valid size), the population proportion of $\mathtt{A}$-players either reaches $x^{\mathsf{N}}=11/28$ and remains there or fluctuates around $x^{\mathsf{N}}=6/7$. However, for larger populations, the fluctuation disappears.
Continuous-time population dynamics. In view of Lemma 2, the continuous-time population dynamics associated with the described discrete population dynamics admit three equilibrium points; one clean-cut equilibrium point $\mathbf{c}^{22}$ at $({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2/28},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}3/28},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0},{0},{0},{3/28},{3/28})$ with the corresponding abstract state $c^{22}$ equals $11/28$, one anticoordinator-driven equilibrium point $\mathbf{a}^{24}$ at $({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2/28},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2/28},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0},{6/28},{8/28},{3/28},{3/28})$ with the corresponding abstract state $a^{24}$ equal to $6/7$, and one coordinator-driven equilibrium point $\mathbf{o}^{23}$ at $({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2/28},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}3/28},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0},{0},{3/28},{3/28},{3/28})$ with the corresponding abstract state $o^{23}$ equal to the threshold of the third coordinating subpopulation, $0.5$. According to Theorem 2, except for a set of initial conditions of measure zero, the population state converges to either of $\mathbf{c}^{22}$ or $\mathbf{a}^{24}$, depending on the initial condition–Figure 4. As for the discrete population dynamics, the observed fluctuations are around ${a}^{24}$. In addition, the equilibrium state at $x^{\mathsf{N}}=11/28$ preserved as the population size grows and is indeed equal to ${c}^{22}$.
Remark 6.
Several studies have analyzed the population dynamics by studying one-dimensional dynamics. In [5], the discrete-time evolution of a heterogeneous population of coordinators was studied by analyzing the evolution of the population proportion of strategy-$\mathtt{A}$ players. A similar approach was used in [19, Section 9] for a heterogeneous population of coordinators and anticoordinators under a synchronous updating rule, where all agents are active at every time step. In [31, 32], scalar continuous-time dynamics instead serve as approximations of the evolution of the population proportion of strategy-$\mathtt{A}$ players in structured populations. In this paper, however, we were interested in studying the evolution of the proportion of strategy-$\mathtt{A}$ players within each subpopulation. Hence, a one-dimensional dynamic would not reveal the behavior of each sub-population. The dimension should be at least as many as the number of sub-populations, i.e., $\mathsf{p}+\mathsf{p}^{\prime}$. Moreover, the evolution of the total proportion of $\mathtt{A}$-players, i.e., $x^{\mathsf{N}}$, is not a Markov chain because the change from $x^{\mathsf{N}}$ to $x^{\mathsf{N}}+1/\mathsf{N}$ (or to $x^{\mathsf{N}}-1/\mathsf{N}$) depends on the proportions of $\mathtt{A}$-players in subpopulations. Nevertheless, the expected change of $x^{\mathsf{N}}$ can be calculated, and that is how we provided the abstract dynamics, which facilitated the subsequent analysis of the $(\mathsf{p}+\mathsf{p}^{\prime})$-dimensional population dynamics.
Figure 4: Example 4. The solution of the continuous-time population dynamics for two different initial conditions. The upper panel shows the evolution of the population state which converges to clean-cut equilibrium point $\mathbf{c}^{22}=({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2,3,0},{0,0,3,3})/28$ and the lower one depicts the evolution of the population state converging to the anticoordinator-driven equilibrium point $\mathbf{a}^{24}=({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2,2,0},{6,8,3,3})/28$, where the corresponding population proportion of $\mathtt{A}$-players equals $\tau_{2}=6/7$. The observed non-smoothness in the evolution of some subpopulations is due to the change in their preferred strategy. For example, in the upper panel, at time zero, the abstract state is below $\tau^{\prime}_{2}$ (resp. $\tau_{3}$) resulting in $\dot{x}^{\prime}_{2}<0$ ($\dot{x}_{3}>0$). As abstract state increases and exceeds $\tau^{\prime}_{2}$ (resp. $\tau_{3}$), the preferred strategy of coordinating (resp. anticoordinating) subpopulation $2$ (resp. $3$) changes to $\mathtt{A}$ (resp. $\mathtt{B}$) yielding $\dot{x}^{\prime}_{2}>0$ (resp. $\dot{x}_{3}<0$).
VII Concluding Remarks
We analyzed the asymptotic behavior of a well-mixed finite heterogeneous population of agents repeatedly choosing between two strategies when the population size approaches infinity. We utilized the available results linking the stationary measures of the Markov chain corresponding to this discrete population dynamics with the steady-state behavior of the associated continuous-time mean dynamics. We then showed that the fluctuations in the finite mixed population of coordinators and anticoordinators likely do not persist with the population size (Figure 2).
The convergence result of the continuous-time population dynamics for exclusive populations of anticoordinators, i.e., 3, matches that of the discrete dynamics [16] where the population state was proven to always reach the clean-cut equilibrium $\mathbf{q}$ or reach the set of two ruffled states (that are not possible here due to the continuity of the state space of the continuous-time population dynamics). Similarly, the convergence result of the continuous-time population dynamics for exclusive populations of coordinators, i.e., 2, matches that of the discrete dynamics [15] where the population state was proven to always equilibrate. Moreover, the equilibria of the discrete and continuous-time dynamics match.
In addition, the stability of clean-cut equilibrium points of continuous-time population dynamics matches that of the discrete dynamics [46]. However, the stability of the ruffled types has not been investigated in the discrete population.
In case of the fluctuations, in all of the examples, we observed that the interval defined by the minimum and maximum of the ratio of strategy $\mathtt{A}$-players in the same population and for a fixed initial condition and different activation sequence contains an anticoordinator-driven equilibrium point of the abstract dynamics. Moreover, as the population size grows, the interval shrinks. It remains yet unknown whether this observation holds true always for every population and whether ruffled anticoordinator-driven equilibria are necessary for the existence of minimal positively invariant sets in the finite discrete populations. The difficulty partly arises from the inherent challenge of characterizing the positively invariant sets and amplitude of the fluctuations. Although [19, Corollary 6] provides bounds on the fluctuations, it still requires the benchmark subpopulations to be known, which appears to be an unsolved complicated problem. The result of this paper holds true in a well-mixed population. Whether a similar result can be obtained for structured populations, such as k-regular networks, is an open question. In this regard, approaches similar to those in [32] might work.
continuous-time population dynamics Stochastic approximation theory Coordinating Anticoordinating Mixture of both $t\ \rightarrow\ \infty$EquilibratesEquilibratesEquilibratesThe fluctuations vanish Population EquilibratesEquilibratesMay Fluctuate Discrete population of finite size $\mathsf{N}$ Coordinating AnticoordinatingMixture of both Discrete population $\mathsf{N}\rightarrow\infty$ Coordinating AnticoordinatingMixture of both $t\ \rightarrow\ \infty$$t\ \rightarrow\ \infty$ Population Population
Figure 2: [Revisit]-The connection between the asymptotic behavior of finite populations of interacting agents evolving based on (3) and their associated continuous-time population dynamics (7) based on Stochastic approximation theory. As population size approaches infinity, the amplitude of the fluctuations in the population proportion of strategy-$\mathtt{A}$ players vanishes with probability $1$.
Appendix-Basic Definitions